OpenStudy (anonymous):

What is the solution to the equation? 32 - 6x + 7x = -15 + 28 x=________

OpenStudy (anonymous):

@paki @agent0smith @hba @mathmate

OpenStudy (welshfella):

first add like term on each side of the equation

OpenStudy (anonymous):

hold up I forgot to add something

OpenStudy (anonymous):

how bout now

OpenStudy (welshfella):

? what did you forget to add?

OpenStudy (anonymous):

X =_________

OpenStudy (welshfella):

oh ok what do you get when you add like terms?

OpenStudy (anonymous):

What do you mean

OpenStudy (welshfella):

well you have like terms -6 + 7x on left and -15 + 28 on the right they can be simplified

OpenStudy (welshfella):

* -6x + 7x on the left

OpenStudy (welshfella):

sorry no i cant do that . CoC does not allow it

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (theedwardsfamily):

@muneeb180 http://www.geteasysolution.com/32-6x+7x=-7+28

OpenStudy (anonymous):

Simplifying 32 + -6x + 7x = -15 + 28 Combine like terms: -6x + 7x = 1x 32 + 1x = -15 + 28 Combine like terms: -15 + 28 = 13 32 + 1x = 13 Solving 32 + 1x = 13 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-32' to each side of the equation. 32 + -32 + 1x = 13 + -32 Combine like terms: 32 + -32 = 0 0 + 1x = 13 + -32 1x = 13 + -32 Combine like terms: 13 + -32 = -19 1x = -19 Divide each side by '1'. x = -19 Simplifying x = -19

OpenStudy (anonymous):

Thank you, @muneeb180